# Tactile SoftHand-A: 3D-Printed, Tactile, Highly-underactuated, Anthropomorphic Robot Hand with an Antagonistic Tendon Mechanism

### v1

@@ 4.3.2 Marker Detection:
We employ the Determinant of Hessian (DoH) estimation method for marker detection to identify white markers in images. The DoH estimation method is a widely-recognized technique in image processing and computer vision for its efficacy in identifying blob features. It operates by locating local maxima of the determinant of the Hessian matrix at every image point. For the given image $I_{gray_{i}}$, the Hessian matrix at a point $(x,y)$ is defined as:
$H(\mathbf{x})=\begin{bmatrix}L_{xx}(\mathbf{x})&L_{xy}(\mathbf{x})\\ L_{yx}(\mathbf{x})&L_{yy}(\mathbf{x})\end{bmatrix},$ (8)
where the components represents the second partial derivatives in the $x$ and/or $y$ directions. We note that a scale parameter $\sigma$ is applied to perform Gaussian Kernel smoothing on the image prior to computing the matrix. This step aids in noise reduction and facilitates feature detection across different scales. Marker detection is achieved by identifying the locations in the image where the Hessian matrix attains a local maximum. The positions of these markers are denoted by $(x_{m},y_{m})$.
@@ 4.3.3 Contact Region Estimation:
Markers within the contact region are influenced by external contacts, levering them towards adjacent uncontacted regions, thereby diminishing the marker density in the region of contact. Consequently, the location of the contact region can be inferred by analyzing the local marker density. Here we adopt an methodology involves the following steps: upon identifying the markers, we compute the marker density using a Gaussian kernel density estimation method. This approach positions Gaussian kernels at the marker centroids, with a constant kernel width $h$ determined by the mean distance between neighboring markers:
@@
where the square-distance norm $\lVert(x,y)-(x_{m},y_{m})\rVert^{2}=(x-x_{m})^{2}+(y-y_{m})^{2}$. We can see from Figure 6(c) that the extent of the low marker density areas in the kernel density map (blue region) depends on the contact depth and reflects the contact area. Finally, we calculate the lowest density point within that region to find the center of contact area ($R_{c}(x_{c},y_{c})$).
$R_{c}(x_{c},y_{c})=\underset{(x,y)\in\mathbb{R}}{\mathrm{arg\,min}}\ \bar{d}(x,y).$ (10)
As shown in Figure 6(d), this represents the contact center (shown as a green circle).
Figure 8: Gestures for the D-type, P-type and A-type fingers under different control inputs to the driving and antagonistic motors (motor inputs 1 to 3, shown to left). The corresponding angles of the MCP, PIP and DIP joints in real time are shown to the right (plots D1-D3 for D-type; P1-P3 for P-type and just A3 for A-type). The A-type finger does not have the ability to move under inputs 1 and 2 due to the absence of soft elements, so there are no plots for A1 and A2. The results in this figure relate to Experiment A.
@@ 5.1 Overall Design
The Tactile SoftHand-A is a major advancement over the previous design of the BPI-SoftHand (Li et al. (2022)), with its use of novel D-type fingers featuring active antagonism, coupled with a modified two-layer palm design for distinct routing of driving and antagonistic tendons, and utilization of two differential mechanisms (Figure 1). We previously introduced three different types of novel fingers; the D-type fingers were selected for their more evenly distributed workspace and used in constructing the Tactile SoftHand-A. Overall, the hand has 15 degrees of freedom and 2 degrees of actuation, with each finger encompassing three joints: MCP, PIP, and DIP, marking an evolution in design and functionality from the previous BPI SoftHand. The overall length of Tactile SoftHand-A is 200 mm with palm width 90 mm and length 26 mm from palm to back cover.
The parts of each finger are depicted in the exploded design illustrations in Figure 2. The hand employs two distinct tendon schemes for the operation of the fingers: one for flexing and another for extending. This dual-tendon system features a soft synergy mechanism, linking each of the five tendons via a spring to its coupling mechanism (tendon spool) and to the differential mechanism, which then connects to a shared spool on the motor, ensuring a coordinated and smooth motion.
@@ 5.2 Differential Mechanism
The Tactile SoftHand-A has two differential mechanisms: one for the active flexing movement of the fingers and one for the active antagonistic movement of the fingers. Each differential mechanism consists of a soft synergy scheme that couples the tendons for flexion and extension separately to springs, with each mechanism connected to a separate tendon coupling driven by a servo motor (see Figure 1). Through the combined differential mechanism, the entire hand can be actively opened and closed with two motors. Hence, the synchronization of all fingers can be achieved easily, and grasping adaptability is also kept, as will be verified in the experimental results.
@@ 5.3 Palm Tendon Layout
To mitigate friction during tendon force transmission, bearing groups were strategically positioned within the palm, as shown in Figure 1(B). The setup for the index finger’s antagonistic tendon illustrates this approach: a bearing set guides the tendon at the finger’s base, preventing direct contact with the base phalanx cover. This arrangement ensures the tendon follows a defined path to the bottom of the palm, where it connects to the guiding mechanism and subsequently to a spring, thus avoiding palm contact and minimizing friction during operation.
@@ 6 Human-guided Control Scheme
When not in contact with an object, using the robotic hand to mirror human hand gestures simplifies control complexity (Gioioso et al. (2013)). Upon contact, the robotic hand can effectively utilize tactile feedback information to automatically adjust its grasping gestures, thereby enhancing its grasping capability. To achieve these objectives, we have developed a tactile feedback control system for Tactile SoftHand-A. It is comprised of four main components (Figure 7): human hand gestures, a PC, a motor controller, and the Tactile SoftHand-A. It operates by translating various hand gestures into control commands. The PC processes raw images of human hands captured by an external camera, along with tactile images from the fingertips of the Tactile SoftHand-A. These images are then converted using the gesture and tactile models into motor control parameters, which are transmitted to the motor controller using UART serial communication. The motor controller primarily utilizes PID control to adjust the rotation of two servo motors, thereby enabling the Tactile SoftHand-A to interact with its surroundings.
@@ 6.1 Open-loop Control Strategy
In experiment A, we use 3 different constant inputs for motors to control P-type, D-type and A-type fingers. By observing the varying gestures of these three types of fingers under the same motor input, we can evaluate the effect of different antagonistic tendon layouts on finger gestures. In Experiments B and C, we implemented an open-loop control strategy without tactile feedback to emphasize the Tactile SoftHand-A’s adaptability to objects’ shape and its capability for varied grasping gestures. Of note, experiment B employed fixed motor inputs to regulate the grasping mechanism of the Tactile SoftHand-A, intentionally reducing the impact of human intervention to underscore its inherent mechanical intelligence. For experiment C, we employed a variety of motor inputs by modifying the encoder difference between the driving and antagonistic motors, demonstrating the Tactile SoftHand-A’s versatility in performing a range of grasping gestures.
@@ 6.2 Closed-loop Control Strategy
In addition, we developed a tactile feedback grasping control system thatintegrates tactile and antagonistic tendon mechanisms to adjust the Tactile SoftHand-A’s grasp gestures via tactile feedback, preventing objects from slipping. Figure 7(A) provides a schematic diagram of the system, while Figure 7(B) illustrates the control flow chart. The system architecture employs a multi-processing approach, as described next. An environmental camera captures real-time images of the human hand, extracts joint landmark points using Mediapipe, and calculates the angles between the fingers and palm. When the tactile sensors in the Tactile SoftHand fingertips do not detect contact, the Tactile SoftHand-A synchronizes with human hand gestures by converting angles into motor inputs through a mapping function. Upon detecting contact, the system switches to tactile feedback closed-loop control, and Tactile SoftHand-A stops synchronizing human hand gestures. The tactile model then checks for slippage by determining if the displacement of the contact point exceeds a threshold. If no slippage is detected, Tactile SoftHand-A maintains the current grasp. If slippage is detected, Tactile SoftHand-A adjusts the grasp gesture and rotates the DIP joints inward to apply greater normal force, preventing the object from slipping. To release an object or grasp a new one, the human hand must be fully opened (with an average angle between the fingers and palm greater than 170 degrees) to exit the tactile feedback closed-loop control. The tactile SoftHand-A will then resynchronize to the human hand gestures until the next contact.
@@ 7 Experiments
7.1 Experiment A - Finger Gestures Comparison
@@
In this finger experiment, we verify the respective controllability of the DIP and PIP joints of the D-type and P-type fingers, by demonstrating how coordinated control of the driving tendon and antagonistic tendon enables active control of the appropriate joints. The A-type finger, due to its antagonistic tendon layout, cannot decouple joints and is used for comparison with other types of fingers. For the D-type, P-type and A-type fingers, Aruco marker codes attached to the phalanges’ sides record the angular changes of the DIP, PIP, and MCP joints, which are then displayed in line graph form (Figure 8). Correspondingly, the input to the motors (encoder values) is also represented in line graphs (Figure 8, motor inputs).
Figure 9: Response of the entire SoftHand-A with D-type fingers to different motor inputs. (A) The three examples of motor input are intended to show that, through the coordinated control of the driving motor and active antagonistic motor, SoftHand-A can achieve changes in grasping gestures through active control of all DIP joints. (B) SoftHand-A gesture changes under these inputs at different points in the trajectory. The red points are taken from the closing trajectory of the SoftHand-A, while the blue points are taken from the opening trajectory of the SoftHand-A. The results in this figure relate to Experiment B.
7.2 Experiment B - Controllable Grasping Gestures
@@
In this hand experiment, we illustrate the capability of the Tactile SoftHand-A to perform controllable grasping gestures. By adjusting the input difference between the driving motor and the antagonistic motor (Figure 9A), we can precisely control the grasping gestures. The behaviour of the hand, displaying both the front and side views under various inputs, is presented in Figure 9B.
7.3 Experiment C - Grasping Adaptability
Figure 10: Grasping adaptability of the tactile SoftHand-A with D-type fingers to various objects with the same motor inputs. (A) Front and side views of the hand at maximal closure. (B) Motor input. (C) Grasp success rate and number of contacting fingertips (validated from the tactile outputs). The grasp success rate is calculated as the number of times an object is successfully grasped in ten grasping experiments. A successful grasp is when the object is grasped for more than 5 seconds under the influence of gravity. The results in this figure relate to Experiment C.
This experiment demonstrates the object-shape adaptation ability of the hand. The Tactile SoftHand-A was mounted on the experimental bench, with fixed motor inputs controlled by the driving and antagonistic motors, enabling the Tactile SoftHand-A to grasp various objects (for the types of motor inputs shown in Figure 10(C)). We selected three objects with constant cross-sections: a hexagonal prism, a triangular prism, and a cylinder; and two multi-section objects: a wine glass and a tendon spool (all objects shown in Figure 10(A)). The single-section objects have a uniform volume and are of height 76 mm, whereas the wine glass and tendon spool measure 136 mm and 89 mm in height, respectively. This setup allowed evaluation of the SoftHand’s adaptability to different objects by comparing the contact conditions of each joint during the grasp. Statistically, we adopted two metrics to assess performance: the average number of contacted fingertips assessed from the tactile sensing, and the proportion of successful grasps (assessed over 10 grasps for each object, varying the placement of the object in hand).
@@ 7.4 Experiment D - Tactile Feedback Control
To evaluate the performance of the tactile feedback control system described (shown in Figure 7), we designed an experiment aimed at demonstrating its capabilities in slip detection and adjusting grasping gestures with tactile feedback. We employed human hands for contact testing, specifically using the thumb tactile sensor of the Tactile SoftHand-A. The experiment is structured into three sequential stages: synchronization, contact and slip. In the synchronization stage, the Tactile SoftHand-A aligns with the gestures of the human hand. During the contact and slip stages, the system engages in closed-loop tactile feedback control. In the slip stage, we assess the system’s ability to detect slip and modify grasping gestures by artificially inducing slip to the thumb of the Tactile SoftHand-A. The results are shown in Figure 13.
Figure 11: Human gesture and tactile control of the Tactile SoftHand-A. (A) human hand gestures used as inputs to control the hand. (B) side view of Tactile SoftHand-A in response to the gesture control. (C) Average degree of human finger closure (1), corresponding motor inputs to the hand (2) and change of finger contact and slip (3). (D) Tactile contact map at various stages throughout the motion as the hand contacts and object. The results in this figure relate to Experiment D.

### v2

@@ 4.3.2 Marker Detection:
We employ the Determinant of Hessian (DoH) estimation method for marker detection to identify white markers in images. The DoH estimation method is a widely-used technique in image processing to identify blob features. For a given image $I_{\rm gray}$, it operates by maximizing the determinant of the Hessian matrix over image points $(x,y)$, using the matrix of second partial derivatives in the $x$ and/or $y$ directions. To reduce noise and aid feature detection across scales, Gaussian kernel smoothing is applied to the image before computing the matrix. Marker detection is achieved by identifying the locations in the image where the Hessian matrix attains local maxima, denoted by $(x_{m},y_{m})$ with $m$ the marker index.
@@ 4.3.3 Contact Region Estimation:
Markers within the contact region are influenced by external contacts, levering them towards adjacent uncontacted regions, thereby diminishing the marker density in the region of contact. Consequently, the location of the contact region can be inferred by analyzing the local marker density. Here we adopt a methodology involving the following steps: upon identifying the markers $(x_{m},y_{m})$, we compute the marker density using a Gaussian kernel density estimation method (Silverman (2018); Lloyd and Lepora (2024); Lu et al. (2025)). This approach positions Gaussian kernels at the marker centroids, with a constant kernel width $h$ taken from the mean distance between neighboring markers:
@@
where the square-distance norm is $\lVert(x,y)-(x_{m},y_{m})\rVert^{2}=(x-x_{m})^{2}+(y-y_{m})^{2}$. We can see from Figure 6(c) that the extent of the low marker density areas in the kernel density map (blue region) depends on the contact depth and represents the contact area. Finally, we calculate the lowest density point within that region:
$(x_{\rm c},y_{\rm c})=\underset{(x,y)\in\mathbb{R}}{\mathrm{arg\,min}}\ \bar{d}(x,y),$ (10)
which, as shown in Figure 6(d), represents the contact center (shown as a green circle).
@@ 5.1 Overall Design
The Tactile SoftHand-A is a major advance over the previous design of the BPI-SoftHand (Li et al. (2022)), with its use of novel D-type fingers featuring active antagonism, coupled with a modified two-layer palm design for distinct routing of agonist and antagonist tendons, and utilization of two differential mechanisms (Figure 1). Earlier in this paper, we introduced three different types of novel fingers; here, we select the D-type fingers for their more evenly distributed workspace to use in constructing the Tactile SoftHand-A. Overall, the hand has 15 degrees of freedom and 2 degrees of actuation, with each finger encompassing three joints: MCP, PIP, and DIP, marking an evolution in design and functionality from the previous BPI SoftHand.
The parts of each finger are depicted in the exploded design illustrations in Figure 2. The hand employs two distinct tendon schemes for the operation of the fingers: one for flexing and another for extending. This dual-tendon system features a soft synergy mechanism, linking each of the five tendons via a spring to its coupling mechanism (tendon spool) and to the differential mechanism, which then connects to a shared spool on the motor, ensuring coordinated and smooth motion.
@@ 5.2 Differential Mechanism
The Tactile SoftHand-A has two differential mechanisms: one for the active flexing movement of the fingers and one for their active extension movement. Each differential mechanism consists of a soft synergy scheme that couples the tendons for flexion and extension separately to springs, with each mechanism connected to a separate tendon coupling driven by a servo motor (see Figure 1). Through the combined differential mechanism, the entire hand can be actively opened and closed with two motors. Hence, the synchronization of all fingers can be achieved easily, and grasping adaptability is also kept, as will be verified in the experimental results.
@@ 5.3 Palm Tendon Layout
To mitigate friction during tendon force transmission, bearing groups were strategically positioned within the palm, as shown in Figure 1(B). The setup for the index finger’s antagonist tendon illustrates this approach: a bearing set guides the tendon at the finger’s base, preventing direct contact with the base phalanx cover. This arrangement ensures the tendon follows a defined path to the bottom of the palm, where it connects to the guiding mechanism and subsequently to a spring, thus avoiding palm contact and minimizing friction during operation.
@@ 6 Human-guided Control Scheme
We developed a comprehensive control strategy for the Tactile SoftHand-A, employing both open-loop and closed-loop control mechanisms. In the closed-loop strategy, the system converts human gestures into motor control signals, utilizing both high-level and low-level controllers (a PC and motor control board). The high-level controller interprets real-time images from the human hand and tactile fingertips, calculating the necessary joint angles, while the low-level controller translates these signals into motor actuation. This dual-level control system allows dynamic adjustments with tactile feedback, ensuring smooth transitions between human gesture synchronization and object anti-slip grasping.
When the robotic hand is not in contact with an object, mimicking human hand gestures simplifies control complexity, as demonstrated in Gioioso et al. (2013). Upon contact, the system switches to closed-loop control, using tactile feedback to automatically adjust the grasping force according to sensor input, preventing slippage. The tactile feedback system, central to the closed-loop control, adapts the hand’s grasping force in response to changes in tactile sensor input. By tracking slippage through marker displacement in tactile images, the system can autonomously adjust the grasp to ensure a secure hold without overexerting force. These adaptive adjustments prevent damage to delicate objects while maintaining control stability.
The control system is an integrated set of modular PID controllers. Specifically, the system consists of four main components (Figure 7): a human-hand gesture detection module, a high-level controller (PC), a low-level controller (motor control board), and the Tactile SoftHand-A hardware. The PC processes raw images of the human hand captured by an environment camera, along with tactile images from the fingertips of the Tactile SoftHand-A. These images, combined with gesture and tactile models, generate motor control parameters transmitted to the motor controller. The motor controller then uses PID control to adjust the rotation of two servo motors, enabling the Tactile SoftHand-A to interact with its environment and adjust its grasp based on real-time tactile feedback.
@@ 6.1 Open-loop Control Strategy
As will be described in more detail in the experiments, we will adopt open-loop feedback control to compare the P-type, D-type, and A-type fingers. In the open-loop strategy, a high-level controller was used to actively adjust the motor inputs of the fingers, while a low-level controller executed these inputs by regulating both the agonist and antagonist tendons. This approach allows us to assess the system’s mechanical intelligence based on its tendon-driven architecture.
The operation involves executing various gestures by setting the desired setpoint on the encoder values of the agonist and antagonist motors. These setpoints range between When the Tactile SoftHand-A is in a fully open gesture and the DIP joint is not actively controlled (values 700 and 820, respectively) and fully closed (200 and 220, respectively). Active control of the DIP joint’s rotation can be achieved under any gesture when the encoder reading of the antagonist motor exceeds that of the agonist motor by approximately 200. When the encoder readings reach approximately 500, the DIP joint achieves its maximum rotational angle.
@@ 6.2 Closed-loop Control Strategy
The closed-loop control strategy integrates tactile and antagonistic tendon mechanisms to adjust the Tactile SoftHand-A’s grasp gestures via tactile feedback, so as to prevent objects from slipping. Figure 7 provides a schematic diagram of the system. The system architecture employs a multi-processing approach, where a camera captures real-time images of the human hand, extracts joint landmark points using the python library MediaPipe, and calculates the angles between the fingers and palm. When the tactile sensors in the SoftHand-A fingertips do not detect contact, the SoftHand-A synchronizes with the human hand gestures by converting angles into motor inputs through a mapping function. Upon detection of contact, the system switches to closed-loop tactile feedback control and the Tactile SoftHand-A stops synchronizing human hand gestures. The tactile model then checks for slippage by determining if the displacement of the contact point exceeds a threshold, similarly to the method considered by Zhang et al. (2024a). If no slippage is detected, the SoftHand-A maintains the current grasp; conversely, if slippage is detected, the SoftHand-A adjusts the grasp gesture and rotates the DIP joints inward to apply a greater normal force, preventing the object from slipping. To release an object or grasp a new one, the human hand must be fully opened (with an average angle between the fingers and palm greater than 170 degrees) to exit the closed-loop tactile feedback control. The tactile SoftHand-A will then resynchronize with the human hand gestures until the next contact. A flow diagram of this process is given in the Appendix (Figure 20).
@@ 7 Experiments
7.1 Experiments A - Evaluation of Single Finger
7.1.1 Experiment A1 - Comparison of Finger Gestures:
@@
This finger experiment investigates the respective controllability and capability to decouple the motion of the DIP and PIP joints of the D-type and P-type fingers, to show how coordinated control of the agonist and antagonist tendons allows active control of the appropriate joints. The A-type finger, which lacks the ability to decouple joints due to its antagonistic tendon arrangement, serves as a comparative model for the other types of fingers. For the D-type, P-type, and A-type fingers, Aruco marker codes affixed to the phalanges’ sides track the angular changes of the DIP, PIP, and MCP joints. The motor inputs (encoder positions) are also taken and depicted as line graphs (Figure 8).
7.1.2 Experiment A2 - Dynamic Performance:
This set of experiments examines the dynamic capabilities of the single fingers. Factors such as response speed, repeatable positioning accuracy, bearing capacity, and manipulation capability are assessed as described below.
A. Response Time Test: The response time of the finger during flexion and extension is evaluated between fully open or closed finger positions. Motor encoder values are used to monitor finger movements and the experiments videoed with a timer for extracting response times.
B. Repeat Positioning Accuracy Test: The finger’s ability to return to predefined positions is assessed with a laser rangefinder. This tracks the fingertip as it moves through 10 distinct positions from fully open to fully closed, with each position repeated over 10 trials.
C. Bearing Capacity Test: The finger’s load-bearing capability is evaluated under static conditions. Weights are applied to the medial phalange until there is greater than 10 mm drop of the fingertip.
D. Finger Manipulation Test: The ability of the entire finger to push a stationary obstacle is assessed by finding the maximum weight moved by extending the finger. This movement is achieved by extending the finger. This experiment acts as a comparison for the next test.
E. Fingertip Manipulation Test: The fingertip’s ability to manipulate objects through controlled movements is assessed by pushing a cuboid inward or outward from its ends and sliding it from above. These fine movements of the fingertip are achieved by coordinating the motion of the MCP and DIP joints. The maximal force exerted is measured.
In all cases, the experimental setup is depicted in the Appendix (Figures 17A-E) and the results summarized in Table 1, with comparison against a single finger of the BPI SoftHand (Li et al. (2022)) that differs primarily in lacking the antagonist tendons and having just one motor.
7.2 Experiments B - Evaluation of Entire SoftHand
7.2.1 Experiment B1 - Controllable Grasping Gestures:
@@
This hand experiment illustrates the capability of the Tactile SoftHand-A to perform controllable grasping gestures. Adjusting the input difference between the agonist and antagonist motors (Figure 9A), can precisely control the grasping gestures, which is assessed by displaying both the front and side views under various inputs.
7.2.2 Experiment B2 - Hand Performance:
This set of experiments assesses the SoftHand-A capabilities for reactive grasping and grasp strength.
A. Response Time Test: The response time of the entire SoftHand during flexion and extension is evaluated between fully open or closed finger positions. The motor encoder values are used to monitor the finger movements, to demonstrate controllable speed and a smooth transition between agonist and antagonist motors over time.
B. Wrench Limit Test: To assess the capability for safe and effective object grasping and manipulation by the SoftHand-A in various application scenarios, we used force sensors to measure the safe wrench limit of the robotic hand under stable grasping conditions.
C. Grasp modes: The SoftHand-A’s capability to perform a power grasp, fingertip grasp and pinch grasp is compared experimentally to highlight the differences between these two modes when gripping objects. In particular, the pinch grasp experiment focuses on the coordination between the thumb and index finger, showcasing the effectiveness of this grasping style for tasks that require precision and control in handling smaller, thinner or more intricate objects. The results are shown in Figure 10.
The experimental setups are depicted in the Appendix (Figure 17A-B) and the results summarized in Table 2, with comparison against the BPI SoftHand (Li et al. (2022)) that differs in lacking the antagonist tendons and their motor.
7.3 Experiment C - Grasping Adaptivity
This experiment demonstrates the object-shape adaptivity of the hand. The Tactile SoftHand-A was mounted on the experimental bench controlled by the agonist and antagonist motors, enabling the Tactile SoftHand-A to grasp various objects (for the motor inputs shown in Figure 12(B)). We selected three regular objects: a triangular prism, a pyramid, and a cylinder; and two irregular objects: a wine glass and a tendon spool. The regular objects are of height 76 mm, whereas the wine glass and tendon spool measure 136 mm and 89 mm in height, respectively.
This setup allowed evaluation of the SoftHand’s adaptivity by comparing the contact conditions of each joint during the various grasps, with two types of grasp per object. In each case the object was placed manually within the hand, which was closed automatically onto the object. To assess performance, we estimated the number of fingertips in contact, assessed from the tactile sensing, These contacts reflect the adaptive grasping ability of the robotic hand.
In addition, we compare the grasping ability of the Tactile SoftHand-A with that of the previous BPI SoftHand that has passive elements. The objects grasped include both YCB objects (Calli et al. (2015)) and common daily items. Grasp success number is calculated as the number of times the object is successfully grasped in 5 grasping experiments. A successful grasp occurs when no change in the contact area is detected within five seconds after the object is grasped under the influence of gravity; i.e., no slippage occurs. The grasping data regarding the original BPI SoftHand is available from a previous study (Li et al. (2022)). The results of both experiments are shown in Figure 12.
@@ 7.4 Experiment D - Tactile Feedback Control
7.4.1 Experiment D1 - Reaction to contact disturbances:
To test the reactivity and sensitivity of the fingertip tactile sensing, we implemented a contact disturbance experiment (see Figure 13). In this experiment, the Tactile SoftHand-A made contact with an object, then the object’s position is continuously disturbed through applying contact force, from which we assess that the hand can maintain contact. We utilize a pen as an indenter to apply arbitrary pressure to the fingertip. The fraction of tactile deformation is estimated by 1 minus the structural similarity index (SSIM) of the indented tactile image from a tactile image of the undeformed state. This tactile deformation was calibrated against normal force in a separate experiment where a single finger is pressed against a force balance. We use a feedback controller that seeks to maintain the tactile deformation to a target setpoint, thereby ensuring the stability of the contact state. The reactivity of the hand can then be assessed from how the finger position moves dynamically in response to the real-time changes in the tactile image.
7.4.2 Experiment D2 - Reactive Teleoperation:
@@
To evaluate the performance of our tactile feedback control system (shown in Figure 7), this experiment uses teleoperation to adjust the grasp gestures using tactile feedback in response to slip. The experiment is structured into three sequential stages: synchronization, contact and slip. In the synchronization stage, the Tactile SoftHand-A aligns with the gestures of the human hand. During the contact and slip stages, the system engages in closed-loop tactile feedback control. In the slip stage, we assess the system’s ability to detect slip and modify grasping gestures by artificially inducing slip to the thumb of the Tactile SoftHand-A, with a probe that we move across the fingertip. The tactile response of the hand uses the closed-loop control strategy described above.
7.4.3 Experiment D3 - Reaction to object disturbances:
The above experiment is then extended to disturbances of objects held in hand. The experiment consists of three stages: synchronization and slip under external disturbance. During the synchronization stage, the BPI SoftHand-A mirrors human hand gestures to actively grasp objects. In the control group, grasping is performed entirely through teleoperation without tactile feedback, and once an object is successfully grasped, the hand maintains its grip. During the slip stage, the system switches to closed-loop tactile feedback control and stops synchronizing with human hand movements. The contact model detects slippage by determining whether the displacement of the contact point exceeds a predefined threshold. If slippage is detected, the Tactile SoftHand-A adjusts its grasp gesture and rotates the DIP joints inward to apply greater normal force, preventing the object from slipping. To quantify the effectiveness of these adaptive grasping strategies, we analyze the variations in normal force recorded by tactile sensors at the fingertips. This force-based evaluation provides quantitative evidence demonstrating how gesture adjustments contribute to improving grasp stability under dynamic conditions.
Figure 8: Experiment A1: Motion of the D-type, P-type and A-type fingers under different control inputs to the agonist and antagonist motors (motor inputs 1 to 3, shown to left). The corresponding angles of the MCP, PIP and DIP joints in real time are shown to the right (plots D1-D3 for D-type; P1-P3 for P-type and just A3 for A-type). The A-type finger does not have the ability to move under inputs 1 and 2 due to the absence of soft elements, so there are no plots for A1 and A2. These plots show how it is possible to decouple the motion of the DIP or PIP joints for the D-type or P-type fingers.
Test SoftHand-A BPI SoftHand
A1. Single finger response time closing from full extension to full flexion 0.39 sec 0.66 sec
A2. Single finger response time opening from full flexion to full extension 0.46 sec 0.32 sec
B. Single finger repeat positioning accuracy Mainly within $\pm 0.3$ $m$ Mainly within $\pm 0.5$ $m$
C. Single finger bearing capacity Drop under 8 N load Drop under 0.9 N load
D. Single finger manipulation: pushing 7 N 2 N
E1. Fingertip manipulation: pushing 3.5 N
E2. Fingertip manipulation: pulling 9 N
E3. Fingertip manipulation: sliding 1.2 N
Table 1: Experiment A2: Comparative single finger test results between the D-type finger of the SoftHand-A and the BPI SoftHand.
Figure 9: Experiment B1: Response of the entire SoftHand-A (D-type fingers) to motor inputs. (A) The three example motor inputs are intended to show that the SoftHand-A can actively control its open/closure and the DIP joints through the coordinated control of the agonist and antagonist motors. (B) SoftHand-A gesture changes under these inputs at different times in the trajectory. Red points indicate the closing trajectory and the blue points show the opening trajectory of the SoftHand-A.
Test SoftHand-A BPI SoftHand
A1. Hand closure response time from full extension to full flexion 0.46 sec 1.38 sec
A2. Hand opening response time from full flexion to full extension 0.59 sec 0.96 sec
B3. Maximum wrench forces (in $x$, $y$, $z$) 20 N, 8 N, 8 N 11.4 N, 4.8 N, 3.5 N
B3. Maximum wrench torques (around $x$, $y$, $z$) 0.6 Nm, 0.3 Nm, 0.6 Nm 0.23 Nm, 0.24 Nm, 0.5 Nm
C. Grasp modes Power, Fingertip, Pinch Power
Table 2: Experiment B2: Comparative whole hand test results between the SoftHand-A and the BPI SoftHand.
Figure 10: Comparison between power grasp (A) fingertip grasp (B) and pinch grasps (C). The difference between the power and fingertip grasp are shown with the two different grasping modes of human hand (a3,b3). Front (a1,b1) and side (a2,b2) views are shown for the power and fingertip grasps. Pinch grasp results are shown with a range of objects held in different ways (c1-c6).